Session: Sep 1997 (session 3)
Frequency: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Onsala (John Conway at Sat Sep 13 16:36:08 METDST 1997) Minor failures
Due to high winds (> 15 m/s) antennas stowed after appoximately 01:30 UT. No other known problems.

Jodrell (Alastair Gunn at Sun Sep 14 13:54:31 METDST 1997) Minor failures
Telescope at wrap limit at start. On source at 16:27 UT. Tape didnt move between 19:50 and 21:15. FS crash between 04:10 and 05:00, started scan 13F at 05:22. FS crash again between 06:53 and 17:03, tape re-positioned for scan 14R at 07:16. Schedules A and B patched together with RW for tape re-positioning. Bad RFI. No other known problems.

Effelsberg (dgraham at Mon Sep 15 10:55:54 METDST 1997) Minor failures
Merged drg files 31a and b to gg031.drg. Source change ref-plsr typically takes 18-20sec. Rain, sometimes heavy.

Torun (K. Borkowski at Wed Sep 17 00:01:58 METDST 1997) Failure
Problems with the recorder. We had to replace the headstack. No data recorded.

Calibration (Fredrik T Rantakyro at Wed Nov 5 15:43:21 MET 1997) Minor failures
The ANTAB formatted TSYS information can be found on VLBEER in the SEP97 directory. The file is called gg031.antab and contains the latest antenna calibration information.